# Integrated Push-and-Pull Update Model for Goal-Oriented Effective Communication

### v1

@@ I Introduction
Messages, e.g., in the form of status update packets, are communicated over existing networked intelligent systems mostly using a push-based communication model. Therein, packets arrived at the source are sent to the destination based on decisions made by the source, regardless of whether or not the endpoint has requested for or plans to utilize these updates to accomplish a goal. In contrast, in a pull-based model, the endpoint decides to trigger and requests packet transmissions from the source and controls the time and the type of generated updates [4, 5, 6, 7, 8, 9, 10]. Nevertheless, this model does not consider the availability of the source to generate updates or the usefulness of those updates. To overcome these limitations, we propose an integrated push-and-pull model that involves both agents/sides in the decision-making process, thereby combining push- and pull-based paradigms in a way that mitigates their drawbacks. In either model, decisions at the source or endpoint could influence the effectiveness of communicated updates. Therefore, we can categorize decision policies into effect-aware and effect-agnostic. Under an effect-aware policy, the source adapts its decisions by taking into account the effects of its communicated packets at the endpoint. Likewise, the endpoint raises queries based on the evolution of the source and the expected importance of pulled updates. Under the effect-agnostic policy, however, decisions are made regardless of their consequent effect on the performance.
@@
In this work, we study a time-slotted end-to-end status update system where a sensing agent observes an information source and communicates updates/observations in the form of packets with an actuation agent. The actuation agent then takes actions based on the successfully received updates as a means to accomplish a subscribed goal at the endpoint. We develop an integrated push-and-pull model, which allows both agents to make decisions based on their local policies or objectives. In particular, a transmission controller at the sensing agent decides to either send or drop update packets according to their potential usefulness at the endpoint. On the other side, a query controller at the actuation agent also determines the time instances around which the actuator should perform actions in the form of raising queries. In that sense, effective updates are those that result in the right impact and actuation at the endpoint. Those queries, however, are not communicated to the sensing agent. Instead, the actuation agent acknowledges the sensing agent of effective updates. With prior knowledge that the effectiveness of updates depends on the actuator’s availability to perform actions, the sensing agent can infer the raised queries using those acknowledgments. A time diagram that shows the processes at both agents is depicted in Fig. 1.
@@
We introduce a metric to measure the effectiveness and the significance of updates and derive a class of optimal policies for each agent that makes effect-aware decisions to maximize the long-term expected effectiveness of update packets communicated to fulfill the goal subject to induced costs. To do so, the agent first needs to estimate the necessary system parameters for making the right decisions. Our analytical and simulation results show that the integrated push-and-pull model comes with a higher energy efficiency compared to the push-based model and better effectiveness performance compared to the pull-based one. Moreover, we observe that utilizing effect-aware policies at both agents significantly improves the effectiveness performance of the system in the majority of the cases with a large gap compared to those of periodic and probabilistic effect-agnostic policies at either or both agents. Accordingly, we demonstrate that the solution to find an optimal effect-aware policy at each agent converges to a threshold-based agent decision framework where the agent can timely decide based on an individual lookup map in hand and threshold boundaries computed to satisfy the goal.
[…]
[scale=1]Media/Time_diagram.eps
@@
Figure 1: A time diagram of processes involving the sensing and actuation agents, illustrating interactions and update communications leading to actions.
@@ I-A Related Works
This paper widely broadens prior work on push-based and (query-) pull-based communications by enabling both agents to make decisions so as to maximize the effectiveness of communicated updates in the system. The pull-based communication model has been widely analyzed but not limited to [4, 5, 6, 7, 8, 9, 10]. In [4], a new metric called effective age of information (EAoI), which comprises the effects of queries and the freshness of updates in the form of the AoI [11, 12, 13], is introduced. Query AoI (QAoI), which is similar to the EAoI, is utilized in [5, 6, 7, 8, 9]. Following the same concept as the QAoI, on-demand AoI is introduced in [10, 14]. Probe (query)-based active fault detection where actuation or monitoring agents adaptively decide to probe sensing agents to detect probable faults at the endpoint is studied in [15]. Most prior work has employed a pull-based communication model and focuses on the freshness and timeliness of information. In this work, we consider multiple information attributes and propose a grade of effectiveness metric to measure the effectiveness of updates, which goes beyond metrics, including AoI, EAoI, QAoI, on-demand AoI, Age of Incorrect Information (AoII) [16], and Value of Information (VoI) [17, 18]. In particular, we focus on the freshness of successfully received updates and the timeliness of performed actions as two attributes of interest through the link level, as well as on the usefulness/significance (semantics) of the updates to fulfill the goal at the source level.
This paper extends our prior work [1], which only considers a pull-based model and an effectiveness metric with two freshness and usefulness attributes. As such, [1] conveys a special form of the decision problem we solve here. In this work, we generalize the problem to a push-and-pull communication model, considering that the sensing and actuation agents individually make decisions and converge to a point where they can transmit updates and raise queries, respectively, which maximize the effectiveness of updates and result in the right impact at the endpoint. Importantly, we assume that the source distribution is not known to the actuation agent, and the sensing agent does not have perfect knowledge of the goal. Therefore, the agents must estimate their required parameters separately. This approach is substantially different from the one in our previous work and other state-of-the-art approaches.
@@ II System Model
We consider an end-to-end communication system in which a sensing agent (SA) sends messages in a time-slotted manner to an actuation agent (AA) as a means to take effective action at the endpoint and satisfy a subscribed goal (see Fig. 2). Specifically, the SA observes a source and generates status update packets in each time slot, and a transmission controller decides whether to transmit that observation or not, following a specific policy. We assume that the source has finite-dimensional realizations and that observation at the $n$-th, $\forall n\in\mathbb{N}$, time slot is assigned a rank of importance $v_{n}$ from a finite set $\mathcal{V}=\{\nu_{i}~{}|~{}i\in\mathcal{I}\}$, with =$\mathcal{I}=\{1,2,...,\lvert\mathcal{V}\rvert\}$, based on its significance or usefulness for satisfying the goal, measured or judged at the source level. [^1] To determine the usefulness of an update, we can use the same metavalue approach proposed in [19, Section III-A]. The elements of $\mathcal{V}$ are independent and identically distributed (i.i.d.) with probability $p_{i}=p_{\nu}(\nu_{i})$ for the $i$-th outcome, where $p_{\nu}(\cdot)$ denotes a given probability mass function (pmf). [^2] A more elaborated model could consider the importance of a realization dependent on the most recently generated update at the SA. This implies that a less important update increases the likelihood of a more significant update occurring later, which can be captured utilizing a learning algorithm.
[…]
[scale=0.6]Media/System_Model.eps
Figure 2: End-to-end status update communication to satisfy a subscribed goal.
@@
The AA is assisted by a query controller that decides to raise queries and pull new updates according to a certain policy. A received packet at the AA has a satisfactory or sufficient impact at the endpoint if that update achieves a minimum effectiveness level subject to the latest query raised and the AA’s availability to act on it. An effective update communication is followed by an acknowledgment of effectiveness (E-ACK) signal sent from the AA to the SA to inform about the effective update communication. We assume all transmissions and E-ACK feedback occur over packet erasure channels (PECs), with $p_{\epsilon}$ and $p_{\epsilon}^{\prime}$ being the erasure probabilities in the forward communication and the acknowledgment links, respectively. Therefore, an E-ACK is not received at the SA due to either ineffective update communication or erasure in the acknowledgment (backward) channel. With this interpretation, channel errors lead to graceful degradation of the proposed scheme. A raised query does not necessarily need to be shared with the SA. As discussed in Section III-C, the SA can deduce the raise of a query or the availability of the AA to take action from a successful E-ACK, given prior knowledge that an update can be effective only if it arrives within the period during which the AA is available to act.
@@
In this model, we consider the goal to be subscribed at the endpoint, with the AA fully aware of it. On the other hand, the SA does not initially know the goal but learns which updates could be useful to accomplish the goal based on the received E-ACK and observations’ significance. Meanwhile, the AA is not aware of the evolution of the source or the likely importance of observations, attempting to approximate it from arrival updates. Consequently, the agents might use different bases to measure the usefulness of the updates and may need to adjust their criteria or valuation frameworks to account for possible changes in goals over time. Finally, we assume that update acquisition, potential communication, and waiting time for receiving an E-ACK occur in one slot.
@@ II-A2 Pull-based
In this model, the query controller plays a central role in the generation of update arrivals at the AA by pulling those updates from the SA. Here, the AA can only take action when queries are raised. However, this model excludes the SA from generating and sending updates.
@@ II-B1 Effect-agnostic
This policy uses a predetermined schedule or random process (e.g., Poisson, binomial, or Markov [5, 6, 7, 8, 9]) to send updates (raise queries) from (by) the SA (AA), without accounting for their impact at the destination. We define a controlled update transmission (query) rate specifying the expected constant number of updates (queries) to be communicated (raised) within a period. Also, as the effect-agnostic policy does not consider what might be happening in the other agent during the time of the decision, there exists an aleatoric uncertainty associated with random updates (queries).
@@ III-A Grade of Effectiveness Metric
We introduce a grade of effectiveness (GoE) metric that comprises several qualitative attributes and characterizes the amount of impact an update makes at the endpoint. Mathematically speaking, the GoE metric is modeled via a composite function $\operatorname{GoE}_{n}=(f\circ g)(\mathcal{I}_{n})$ for the $n$-th time slot. Here, $g:\mathbb{R}^{x}\to\mathbb{R}^{y},x\geq y$, is a (nonlinear) function of $x\in\mathbb{Z}^{+}$ information attributes $\mathcal{I}_{n}\in\mathbb{R}^{x}$, and $f:\mathbb{R}^{y}\to\mathbb{R}$ is a context-aware function. [^3] The GoE metric in this form can be seen as a special case of the SoI metric introduced in [2, 20, 19]. The particular forms of functions $f$ and $g$ could vary according to different subscribed goals and their relevant requirements.
In this paper, without loss of generality, we consider freshness of updates and timeliness of actions as the main contextual attributes. The first comes in the form of age of information (AoI) metric, which is denoted by $\Delta_{n}$. The second is measured from the action’s lateness, denoted by $\Theta_{n}$. Thereby, we can formulate the GoE metric as follows
$\operatorname{GoE}_{n}=f_{g}(g_{\Delta}(\hat{v}_{n},\Delta_{n}),g_{\Theta}(\Theta_{n});g_{c}(C_{n}))$ (1)
where $C_{n}$ represents the overall cost incurred in the $n$-th time slot. Also, $g_{\Delta}:{\mathbb{R}_{0}^{+}}\times{\mathbb{R}^{+}}\rightarrow\mathbb{R}_{0}^{+}$, $g_{\Theta}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}_{0}^{+}$, and $g_{c}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}_{0}^{+}$ are penalty functions, and $f_{g}:{\mathbb{R}_{0}^{+}}\times{\mathbb{R}_{0}^{+}}\times{\mathbb{R}_{0}^{+}}\rightarrow\mathbb{R}_{0}^{+}$ is a non-decreasing utility function. Moreover, $g_{\Delta}$, $g_{\Theta}$, and $g_{c}$ are non-increasing with respect to (w.r.t.) $\Delta_{n}$, $\Theta_{n}$, and $C_{n}$, respectively, while $g_{\Delta}$ is non-decreasing w.r.t. $\hat{v}_{n}$. Here, $\hat{v}_{n}$ is the usefulness of the received update from the endpoint’s viewpoint in the $n$-th slot. Thus, we assume that $\hat{v}_{n}$ belongs to the set $\hat{\mathcal{V}}=\{0\}\cup\{\hat{\nu}_{j}~{}|~{}\hat{\nu}_{j}>0,j\in\mathcal{J}\}$ with i.i.d. elements, where $\mathcal{J}=\{1,2,...,\lvert\hat{\mathcal{V}}\rvert-1\}$, the $j$-th element has probability $q_{j}=p_{\hat{\nu}}(\hat{\nu}_{j})$, and $p_{\hat{\nu}}(\cdot)$ is a pmf derived in Section V-A. Since the packet is sent over a PEC, $\hat{v}_{n}=0$ if it is erased or the update ends up being useless at the endpoint.
@@ III-A1 AoI
$u(n)=\operatorname{max}\big{\{}m~{}|~{}m\leq n,\beta_{m}(1-\epsilon_{m})=1\big{\}}$ (2)
with $\epsilon_{m}\in\{0,1\}$ being the channel erasure in the $m$-th slot. In addition, $\beta_{m}\in\{0,1\}$ indicates the query controller’s decision, where $\beta_{m}=1$ means pulling the update; otherwise, $\beta_{m}=0$.

### v2

@@ I Introduction
Messages, e.g., in the form of status update packets, are communicated over existing networked intelligent systems mostly using a push-based communication model. Therein, packets arrived at the source are sent to the destination based on decisions made by the source, regardless of whether or not the endpoint has requested or plans to utilize these updates to accomplish a goal. In contrast, in a pull-based model, the endpoint decides to trigger and requests packet transmissions from the source and controls the time and the type of generated updates [4, 5, 6, 7, 8, 9, 10]. Nevertheless, this model does not consider the availability of the source to generate updates or the usefulness of those updates. To overcome these limitations, we propose an integrated push-and-pull model that involves both agents/sides in the decision-making process, thereby combining push- and pull-based paradigms in a way that mitigates their drawbacks. In either model, decisions at the source or endpoint could influence the effectiveness of communicated updates. Therefore, we can categorize decision policies into effect-aware and effect-agnostic. Under an effect-aware policy, the source adapts its decisions by taking into account the effects of its communicated packets at the endpoint. Likewise, the endpoint initiates queries based on the evolution of the source and the expected importance of pulled updates. Under the effect-agnostic policy, however, decisions are made regardless of their consequent effect on the performance.
We illustrate the motivation behind the effect-aware integrated model with a real-world example: Consider a remote patient monitoring system. A medical expert (the endpoint) is responsible for responding to a patient’s vital signs (the source) but may not always be available due to other responsibilities. The push-based model risks overwhelming the expert with irrelevant or redundant information, or worse, causing critical updates to be missed amid competing demands. Conversely, the pull-based model requires the expert to check the patient’s status at predefined intervals, but this rigidity may lead to missing urgent or critical changes in the patient’s condition. The push-and-pull model strikes a balance between these approaches by integrating the strengths of both. It allows the experts to receive timely updates and act while maintaining the flexibility to manage other responsibilities or to respond effectively without dedicating all their time to a single patient. By integrating effect-aware decision-making, updates are prioritized based on their importance and the expert’s availability to act. Furthermore, by having access to and continually monitoring the patient’s evolving health condition, the expert can dynamically adjust query timing and align decisions with the expected relevance of updates.
@@
In this work, we investigate a time-slotted end-to-end status update system where a sensing agent observes an information source and communicates updates/observations in the form of packets with an actuation agent. The actuation agent then takes action based on the updates successfully received as a means to accomplish a subscribed goal at the endpoint. We develop an integrated push-and-pull model, which allows both agents to make decisions based on their local policies or objectives. In particular, a transmission controller at the sensing agent decides to either send or drop update packets according to their potential usefulness at the endpoint. On the other side, a query controller at the actuation agent also determines the time instances around which the actuator should perform actions in the form of raising queries. In that sense, effective updates are those that result in the right impact and actuation at the endpoint. Those queries, however, are not communicated to the sensing agent. Instead, the actuation agent acknowledges the sensing agent of effective updates. With prior knowledge that the effectiveness of updates depends on the actuator’s availability to perform actions, the sensing agent can infer the initiated queries using those acknowledgments. A time diagram showing processes at both agents is depicted in Fig. 1.
@@
We introduce a metric to measure the effectiveness and the significance of updates and derive a class of optimal policies for each agent that makes effect-aware decisions to maximize the long-term expected effectiveness of update packets communicated to fulfill the goal subject to induced costs. These costs may arise from transmitting updates, initiating queries, and the availability of the actuation agent to respond and take action, all of which depend on the communication model. To do so, the agent first needs to estimate the necessary system parameters for making the right decisions. Our analytical and simulation results show that the integrated push-and-pull model comes with a higher energy efficiency compared to the push-based model and better effectiveness performance compared to the pull-based one. Moreover, we observe that utilizing effect-aware policies at both agents significantly improves the effectiveness performance of the system in the majority of the cases, with a large gap compared to periodic and probabilistic effect-agnostic policies at either or both agents. Accordingly, we demonstrate that the solution to find an optimal effect-aware policy at each agent converges to a threshold-based agent decision framework where the agent can timely decide based on an individual lookup map in hand and threshold boundaries computed to satisfy the goal.
@@
Figure 1: A timing diagram of processes involving the sensing and actuation agents, illustrating interactions and update communications leading to actions.
@@ I-A Related Works
This paper widely broadens prior work on push-based and (query-) pull-based communications by enabling both agents to make decisions so as to maximize the effectiveness of communicated updates in the system. The pull-based communication model has been widely analyzed but not limited to [4, 5, 6, 7, 8, 9, 10]. In [4], a metric called effective age of information (EAoI), which comprises the effects of queries and the freshness of updates in the form of the age of information (AoI) [11, 12, 13], is introduced. Query AoI (QAoI), which is similar to the EAoI, is utilized in [5, 6, 7, 8, 9]. Following the same concept as the QAoI, on-demand AoI is introduced in [10, 14]. Probe (query)-based active fault detection where actuation or monitoring agents adaptively decide to probe sensing agents to detect probable faults at the endpoint is studied in [15]. Most prior work has employed a pull-based communication model and focuses on the freshness and timeliness of information. In this work, we consider multiple information attributes and propose a grade of effectiveness metric to measure the effectiveness of updates, which goes beyond metrics, including AoI, EAoI, QAoI, on-demand AoI, age of incorrect information (AoII) [16], and value of information (VoI) [17, 18]. In particular, we focus on the freshness of successfully received updates and the timeliness of performed actions as two attributes of interest through the link level, as well as on the usefulness/significance (semantics) of the updates to fulfill the goal at the source level.
This paper extends our prior work [1], which only considers a pull-based model and an effectiveness metric with two freshness and usefulness attributes. As such, [1] conveys a special form of the decision problem we solve here. In this work, we generalize the problem to a push-and-pull communication model, considering that the sensing and actuation agents individually make decisions and converge to a point where they can transmit updates and initiate queries, respectively, which maximize the effectiveness of updates and result in the right impact at the endpoint. Importantly, we assume that the source distribution is not known to the actuation agent, and the sensing agent does not have perfect knowledge of the goal. Therefore, the agents must estimate their required parameters separately. This approach is substantially different from the one in our previous work and other state-of-the-art approaches.
@@ II System Model
We consider an end-to-end communication system in which a sensing agent (SA) sends messages in a time-slotted manner to an actuation agent (AA) as a means to take effective action at the endpoint and satisfy a subscribed goal. Specifically, the SA observes a source and generates status update packets in each time slot, and a transmission controller decides whether to transmit that observation or not, following a specific policy. We assume that the source has finite-dimensional realizations and that observation at the $n$-th, $\forall n\in\mathbb{N}$, time slot is assigned a rank of importance $v_{n}$ from a finite set $\mathcal{V}=\{\nu_{i}\,\lvert\,i\in\mathcal{I}\}$, with =$\mathcal{I}=\{1,2,\dots,\lvert\mathcal{V}\rvert\}$, based on its significance or usefulness for satisfying the goal, measured or judged at the source level. [^1] To determine the usefulness of an update, we can use the same metavalue approach proposed in [19, Section III-A]. The elements of $\mathcal{V}$ are independent and identically distributed (i.i.d.) with probability $p_{i}=p_{\nu}(\nu_{i})$ for the $i$-th outcome, where $p_{\nu}(\cdot)$ denotes a given probability mass function (pmf). [^2] A more elaborated model could consider the importance of a realization dependent on the most recently generated update at the SA. This implies that a less important update increases the likelihood of a more significant update occurring later, which can be captured utilizing a learning algorithm.
@@
The AA is assisted by a query controller that decides to initiate queries and pull new updates according to a certain policy. A received packet at the AA has a satisfactory or sufficient impact at the endpoint if that update achieves a minimum effectiveness level subject to the latest query initiated and the AA’s availability to act on it. An effective update communication is followed by an acknowledgment of effectiveness (E-ACK) signal sent from the AA to the SA to inform about the effective update communication. We assume all transmissions and E-ACK feedback occur over packet erasure channels (PECs), with $p_{\epsilon}$ and $p_{\epsilon}^{\prime}$ being the erasure probabilities in the forward communication and the acknowledgment links, respectively. Therefore, an E-ACK is not received at the SA due to either ineffective update communication or erasure in the acknowledgment (backward) channel. With this interpretation, channel errors lead to graceful degradation of the proposed scheme. An initiated query does not necessarily need to be shared with the SA. As discussed in Section III-C, the SA can deduce the initiation of a query or the availability of the AA to take action from a successful E-ACK, given prior knowledge that an update can be effective only if it arrives within the period during which the AA is available to act.
@@
In this model, we consider the goal to be subscribed at the endpoint, with the AA fully aware of it. On the other hand, the SA does not initially know the goal but learns which updates could be useful to accomplish the goal based on the received E-ACK and observations’ significance. Meanwhile, the AA is not aware of the evolution of the source or the likely importance of observations, attempting to approximate it from arrival updates. Consequently, the agents might use different bases to measure the usefulness of the updates and may need to adjust their criteria or valuation frameworks to account for possible changes in goals over time. Finally, we assume that update acquisition, potential communication, and waiting time for receiving an E-ACK occur within one slot.
@@ II-A2 Pull-based
In this model, the query controller plays a central role in the generation of update arrivals at the AA by pulling those updates from the SA. Here, the AA can only take action when queries are initiated. However, this model excludes the SA from generating and sending updates.
@@ II-B1 Effect-agnostic
This policy employs a predetermined schedule or random process (e.g., Poisson, binomial, or Markov [5, 6, 7, 8, 9]) to send updates (initiate queries) from (by) the SA (AA), without considering their impact at the endpoint. We define a controlled update transmission (query) rate specifying the expected constant number of updates (queries) to be communicated (initiated) within a period. Also, as the effect-agnostic policy does not account for what might be happening in the other agent during the decision time, there exists an aleatoric uncertainty associated with random updates (queries).
@@ III-A Grade of Effectiveness Metric
We introduce a grade of effectiveness (GoE) metric that comprises several qualitative attributes and characterizes the amount of impact an update makes at the endpoint. Mathematically speaking, the GoE metric is modeled via a composite function $\operatorname{GoE}_{n}=(f\circ g)(\mathcal{I}_{n})$ for the $n$-th time slot. Here, $g:\mathbb{R}^{x}\to\mathbb{R}^{y},x\geq y$, is a (nonlinear) function of $x\in\mathbb{N}$ information attributes $\mathcal{I}_{n}\in\mathbb{R}^{x}$, and $f:\mathbb{R}^{y}\to\mathbb{R}$ is a context-aware function. [^3] The GoE metric in this form can be seen as a special case of the semantics of information (SoI) metric introduced in [2, 20, 19]. The particular forms of functions $f$ and $g$ could vary according to different subscribed goals and their relevant requirements.
In this paper, without loss of generality, we consider freshness of updates and timeliness of actions as the main contextual attributes. The first comes in the form of the AoI metric, which is denoted by $\Delta_{n}\in\mathbb{N}$. The second is measured from the action’s lateness, denoted by $\Theta_{n}\in\mathbb{N}$. Thereby, we can formulate the GoE metric as follows
$\operatorname{GoE}_{n}=f_{g}\Big{(}g_{\Delta}(\hat{v}_{n},\Delta_{n}),g_{\Theta}(\Theta_{n});g_{c}(C_{n})\Big{)}$ (1)
where $C_{n}\in\mathbb{R}_{0}^{+}$ represents the overall cost incurred at the $n$-th time slot. Also, $g_{\Delta}:\hat{\mathcal{V}}\times\mathbb{N}\rightarrow\mathbb{R}_{0}^{+}$, $g_{\Theta}:\mathbb{N}\rightarrow\mathbb{R}_{0}^{+}$, and $g_{c}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}_{0}^{+}$ are penalty functions, and $f_{g}:{\mathbb{R}_{0}^{+}}\times{\mathbb{R}_{0}^{+}}\times{\mathbb{R}_{0}^{+}}\rightarrow\mathbb{R}_{0}^{+}$ is a non-decreasing utility function. Moreover, $g_{\Delta}$, $g_{\Theta}$, and $g_{c}$ are non-increasing with respect to (w.r.t.) $\Delta_{n}$, $\Theta_{n}$, and $C_{n}$, respectively, while $g_{\Delta}$ is non-decreasing w.r.t. $\hat{v}_{n}$. Here, $\hat{v}_{n}$ is the usefulness of the received update from the endpoint’s viewpoint at the $n$-th slot. Thus, we assume that $\hat{v}_{n}$ belongs to the set $\hat{\mathcal{V}}=\{0\}\cup\{\hat{\nu}_{j}\,\lvert\,\hat{\nu}_{j}>0,j\in\mathcal{J}\}$ with i.i.d. elements, where $\mathcal{J}=\{1,2,\dots,\lvert\hat{\mathcal{V}}\rvert-1\}$, the $j$-th element has probability $q_{j}=p_{\hat{\nu}}(\hat{\nu}_{j})$, and $p_{\hat{\nu}}(\cdot)$ is a pmf derived in Section V-A. Since the packet is sent over a PEC, $\hat{v}_{n}=0$ if it is erased or the update ends up being useless at the endpoint.
@@ III-A1 AoI
$u(n)=\operatorname{max}\big{\{}m\,\lvert\,m\leq n,\beta_{m}(1-\epsilon_{m})=1\big{\}}$ (2)
with $\epsilon_{m}\in\{0,1\}$ being the channel erasure at the $m$-th slot. In addition, $\beta_{m}\in\{0,1\}$ indicates the query controller’s decision, where $\beta_{m}=1$ means pulling the update; otherwise, $\beta_{m}=0$.